OpenStudy (loser66):

Explain me, please. I know how to do it but I want to know the logic on it rather than just apply formula. Problem, find all solutions of 3x + 4y = 7

OpenStudy (loser66):

gcd (3, 4 ) =1 4 = 3 + 1 1 = 4 - 3 --> 1 = 3 (-1) + 4(1) multiple all by 7 , we get \(3 (-7) +4(7) =7\) Hence \(x_0= -7 \\y_0 = 7\) \(x = -7 \huge \color{red}{+}\)4t \(y = 7 \huge \color{blue}{-}\) 3t I would like to know the logic on those \(\color{red}{+}\) and \(\color{blue}{-}\)

OpenStudy (loser66):

@ganeshie8

OpenStudy (anonymous):

You can find the proof for why the solution of a linear diophantine solution has that particular form here: http://math.uh.edu/~minru/5333-07/hj02cs01.pdf